OpenStudy (anonymous):

The formula for projectile motion is: h(t) = -16t2 + vot + so, where h is the height (at a given time t), vo is the initial velocity and so is the starting height. A rocket is launched straight up from the top of an 64 foot building with the initial velocity of 160 feet per second. Find the time interval when the rockets' height exceeds 400 feet.

ganeshie8 (ganeshie8):

where are you stuck

OpenStudy (anonymous):

how do i solve this??

ganeshie8 (ganeshie8):

we start by interpreting the question and figure out what stuff we're given, and what exactly we're asked to solve.

OpenStudy (imstuck):

you can fill into your equation the values you are given. Ooops, I won't interfere with ganeshie8's lesson...I dislike very much when people do that to me! Good luck with this! You're in good hands...

ganeshie8 (ganeshie8):

read the question and can you tell me below : 1) what we're given ? 2) what we need to solve ?

OpenStudy (anonymous):

given is the height of the building and the original velocity

ganeshie8 (ganeshie8):

thats right ! and what are we trying to solve ?

OpenStudy (anonymous):

the time interval when the rockets' height exceeds 400 feet.

ganeshie8 (ganeshie8):

Excellent!! so we're given below : initial height, \(S_0 = 64 \) feet initial velocity, \(V_0 = 160\) feet per second and we need to find \(\large t\) at which the height equals \(400\) feet

ganeshie8 (ganeshie8):

plugin the values in the given equation for height

ganeshie8 (ganeshie8):

\[\large h(t) = -16t^2 + V_0 t + S_0,\]

OpenStudy (anonymous):

t is 2.5?

ganeshie8 (ganeshie8):

\[\large 400 = -16t^2 + 160 t + 64\]

ganeshie8 (ganeshie8):

nope, try again. we need to solve above quadratic ^

OpenStudy (anonymous):

you minus 64 from both sides right

ganeshie8 (ganeshie8):

im thinking of dividing 16 both sides first

ganeshie8 (ganeshie8):

\[\large \dfrac{400}{16} = \dfrac{-16t^2 + 160 t + 64}{16}\]

OpenStudy (anonymous):

25=160t+64

ganeshie8 (ganeshie8):

you need to divide each term on the right hand side cuz all are connected by addition..

ganeshie8 (ganeshie8):

\[\large 25 = -t^2 + 10 t + 4\]

ganeshie8 (ganeshie8):

send the right hand stuff to left : \[\large t^2 - 10t +21 = 0\]

ganeshie8 (ganeshie8):

see if you can factor this quadratic

OpenStudy (anonymous):

ok t= 7,3

ganeshie8 (ganeshie8):

Yes !! which one is small ? 7 or 3 ?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

so that goes first 3 then 7

ganeshie8 (ganeshie8):

so, at t = 3 the rocket's height reaches 400 feet

OpenStudy (anonymous):

ok it wants the interval notation so its (3,7)?

ganeshie8 (ganeshie8):

that looks good ^

OpenStudy (anonymous):

ok good!!!

ganeshie8 (ganeshie8):

between t = 3 and t=7, the rocket's height stays above 400

OpenStudy (anonymous):

ok cool

ganeshie8 (ganeshie8):

good job !!

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

ok how about something like this?

OpenStudy (anonymous):

The equation for the calculation of the parallel resistance R1 and R2 is given by: Rtotal = R1R2 R1 + R2 If the parallel circuit R2 = 4 ohms, find the interval in which Rtotal is at most 2 ohms.

OpenStudy (anonymous):

R1R2 over R1+R2

ganeshie8 (ganeshie8):

\[\large R_{\text{total}} = \dfrac{R_1 R_2}{R_1 + R_2}\]

ganeshie8 (ganeshie8):

like that ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

Alright, so what stuff are we given, and what we need to solve exactly ?

OpenStudy (anonymous):

R2 = 4 ohms

OpenStudy (anonymous):

R TOTAL is at the most 2 ohms

ganeshie8 (ganeshie8):

that means Rtotal cannot exceed 2ohms

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so it would either be [0, 4]

ganeshie8 (ganeshie8):

lets see

OpenStudy (anonymous):

you would divide right?

OpenStudy (anonymous):

whats that sign mean? after the 4?

ganeshie8 (ganeshie8):

wait a sec, actually you're correct ! i made a mistake... sorry :/

OpenStudy (anonymous):

ok so its 0,4

ganeshie8 (ganeshie8):

\[\large 2 \ge \dfrac{4R_1 }{R_1 + 4} \] \[\large 2R_1 + 8 \ge 4R_1 \] \[\large R_1 \le 4 \]

ganeshie8 (ganeshie8):

\(\large [0, 4]\) is right !

OpenStudy (anonymous):

ok cool thanks

OpenStudy (anonymous):

can we do another?

OpenStudy (anonymous):

Sue drove her car at 25 mph for the first part of the trip and then changed her speed to 75 mph for the second part. If the entire trip is 725 miles and took 13 hours, how many hours are traveled at each speed?

ganeshie8 (ganeshie8):

hey please post it as a new question...

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

ty :) when possible, please limit to one question per thread :)

OpenStudy (anonymous):

ok ty:)